First City Bank pays 8 percent simple interest on its savings account balances, whereas Second City Bank pays 8 percent interest compounded annually. 


If you made a $61,000 deposit in each bank, how much more money would you earn from your Second City Bank account at the end of 8 years? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) 


Difference in accounts =        
Answer #1

First city bank:

Simple interest per year = Principal * Rate = 61000 * 8% = 4880

Interest earned for 8 years = 4880 *8 = 39040

Balance at end of 8 years = 61000 + 39040 = 100040

Second city bank:

Future value = Principal * (1+R)^N

= 61000 * (1+8%)^8

= 112906.74

Excess interest earned from investment in second city bank = 112906.74 - 100040 = 12866.74
Answer #2

result = 2632.63

answered by: eloi.yarar.19@first